Solve the system of equation.
-x + 2y = 1
0.875x + 2y = 2.875

Answer by checkley77(12844)

-x + 2y = 1
0.875x + 2y = 2.875 SUBTRACT
-------------------------------
-1.875X=-1.875
X=-1.875/-1.875
X=1 ANS.
-1+2Y=1
2Y=1+1
2Y=2
Y=2/2
Y=1 ANS.
PROOF:
.875*1+2*1=2.875
.875+2=2.875
3.875=2.875